A poll conducted in 2017 found that 52% of U.S. adult Twitter users get at least some news on Twitter. The standard error for this estimate was 2.4%, and a normal distribution may be used to model the sample proportion. Construct a 99% confidence interval for the fraction of U.S. adult Twitter users who get some news on Twitter, and interpret the confidence interval in context.
Answer #1

standard error, SE = 0.024
pcap = 0.52

Given CI level is 99%, hence α = 1 - 0.99 = 0.01
α/2 = 0.01/2 = 0.005, Zc = Z(α/2) = 2.58

CI = (pcap - z*SE , pcap + z*SE)
= (0.52 - 2.58*0.024 , 0.52 + 2.58*0.024)
= (0.4581, 0.5819)

The calculated CI indicates that the proportion of U.S. adult Twitter users who get some news on Twitter lies within (0.4581, 0.5819) with 99% confidence.